Restart Night #9 UT 30 Aug 2020

OSA: S. Allanson

ISp: G. Bechetii

Mtn Manager: B. Bingham

ISA: O. Kuhn


We were closed all night due to clouds and rain. The crane troughs were deployed after which the telescope had to stay at zenith, so this precluded any binocular LUCI flexure tests. Dave worked on the LUCI2 wavelength calibration and Greg on AO checkout using LUCI1. I ran some closed dome checks on LBCB: biases and pseudo-flats in all filters. LBCB sci and tracker images look normal and the pseudo-flats show no signs of debris. The LBCR pseudo-flats show the squiggle in the upper right corner.

MODS1B is back in business after the microLynx replacement and I took some MODS comps and also some long hatch-closed exposures to check for IR laser spots; these were quite obvious in a stacked sky-subtracted 2D spectrum taken a few nights ago.


There were some issues:
  • LBCR dewar going soft. IT 8177
  • LUCI2 FPU temperature gradually rising. No IT on this one yet.
  • The ds9 launched by RB_Science would disappear/reappear so that, for biases and high cadence observations, it was not possible to view the data real-time. IT 8176

Proposed Plan:

  • PFS Alignment Alignment - daytime - into handover if needed - CRITICAL
  • Focus check time permitting (or if weather keeps us closed at sunset)
  • L! Rotator Center (0.5 hr) - high
    • M92, M34, or M52 - fields available for all night
  • L1 Rotator ZP for AGW - (1 hr) - critical
    • Long Slit Method - instructions to be provided by J. Power
      • Suggest doing 3 "masks" to verify the RotZP (based on experience with issues from L2 run)
  • AGW Transform for both LUCIs (use the N30 fieldstop, H or FeII filters) - (2 hr) - high/critical
  • LUCI-1 Imaging check (can run LUCI-2 in parallel to repeat) - high
    • Important to test in and out of field to "stress-test" AGW-1 - does it recover from long throws? and multi-point dithers?
    • Sidereal
      • Dither in-field
        • J1601 (sets around 10pm) - dithering in the field -Kband only
        • NGC 7027 (sets 4am) - cycles through all filters, can stop early
      • Dither out-of-field
        • M52 - (all night) - KJH
        • NGC 40 - (all night) - all filters
    • NS
      • Comet 29/P - (after 10pm) - JHK - need to update for tonight's ephemeris
  • Twilight Flats - low
  • Photometric Standards - (1-2hrs) - medium - only if clear - any seeing
  • AFTER LUCI-1 checkout:
  • LBCs
    • Collimate WRS/DOFPIA - high (4 x 0.25 hrs) - to check LBC-Blue
    • Superfocus sequence -PA=0,180 high (1 hr) - to check LBC-Blue
    • Rotator Center Checks, start with PA=0, 180 (track mode), then 90,270 - high (1hr) - to check LBC-Blue
    • Test Science OB - guiding - medium (1hr)
      • Sidereal
      • Non-Sidereal
    • Standard Photometric fields - requires clear skies - any seeing
    • Spectrophotometric Standards - requires clear skies - any seeing
    • Twlight Flats
  • MODS
    • Spectrophotometric stars if clear (and if not clear only if taking data)
    • Photometric Standards - (1 hr) - only if clear - any seeing
    • Mixed Mode LBC-R/MODS-2
    • Skyflats MODS2-Blue
    • Revisit Transient - (1hr) - low-medium priority - not critical for redshift, if initial spectra look interesting this may be useful

Closed Dome Activities

  • LUCI-Bino Flexure
  • LUCI-Bino Wavelength Calibration
  • LBC-B checkout (pseudo dome flats and biases)
  • MODS-1Blue - mechanical checkout after microlynx swap
  • AO? not compatible with LUCI-Bino


UT 01:57 There are clouds and predictions for heavy rain.

UT 2 Dave is online, doing the LUCI wavelength calibration.

UT 03:55 It is raining steadily up there. Bruce and Steve are deploying the crane troughs. There are no drips in the chamber yet but as the forecast is for heavy rain.

UT 04:30 Around this time, Greg wrapped up (a thunderstorm took out his internet). Dave is continuing with the voltage-wavelength calibrations.

UT 06:12 I am waking both MODS and took a sieveSnap. 20200830.0001. All 4 sieveSnaps look good (The grating unit of MODS1 Blue is again working).

UT 06:22 Dave is signing off - he obtained wavelength calibration data, but as the trough is deployed we cannot move off zenith and this pretty much precludes the collection of flexure data. (We could only do one elevation.) He noted, however, that the FPU temperature has been steadily rising since about 4pm. The temperature has rising 7.5 deg C in the past 7 hours. It may be heat load on the mask in the FPU, so Dave put the mask in storage just now.

UT 06:40 Taking 25 LBC biases, 2x. Got an RPC error on the red side.
  • RB_Science ds9 keeps disappearing and popping up anew, so quickly that almost no images are actually displayed to it. Filed IT 8176
  • RPC error with LBCR. lbckill, lbcstart done and a 3rd set of biases has been started. This action appears to have resolved the error.
2020/08/30 07:03:45.147396 W R CAMERA            clnt_create failed: src/camera/camera.c: RPC: Remote system error - Connection refused  [src/camera/camera.c:1143]
2020/08/30 07:03:45.147453 E R CAMERA            RPC Client/Host communication failed (4) [src/camera/camera.c:1169]
2020/08/30 07:15:26.473340 N -                   UI update with message >(RED Channel) OB execution error: (CAMERA) RPC Client/Host communication failed!< at 1598771726.473143

UT 06:40 Taking MODS comp lamp calibrations, these will be mods#c.20200830.00NN.fits.
  • dual
  • red-only
UT 07:40 Take closed-hatch spectra with and without the IR laser enabled to check on the "IR laser spots"
  • hatch closed, calib out, agw homed
  • 0.6" slit
  • dual grating
  • irlaser enabled: m1b 5, m1r 7, m2b 5, m2r 8
  • irlaser disabled: m1b 6, m1r 8, m2b 6, m2r 9
  • the difference image m2r shows lots of streaks on the lower right quadrant - seen before.
UT 08:18 LUCI2 FPU temperature: Moving the mask caused a drop in temperature, but only for a short time. It started then to rise again, pretty much continuing the curve from earlier. It is now at 91.18 K.

UT 09:25 ALH went off --- LBCR dewar pressure is high, at 1.01-e04 mbar. There has been a gradual increase in dewar pressure since Aug 23. IT 8177 filed.

UT 11:08 LBCB+LBCR: I am running the OB to cycling through all filters with lower floods on. (I guess I did not click "play" until 12:00, though...) lbcb.20200830.12* and lbcr.20200830.12*.The LBCR flats do show the "squiggle" at the top right of the FOV.

UT 12:00 LUCI2 FPU temperature is now up to 92.7 K.

UT 13:00 LUCI2 FPU to 93

UT 13:15 Completed another irlaser disabled/enabled pair with MODS.
  • irlaser disabled: m1b7, m1r 9, m2b 7, m2r 10
  • irlaser enabled: m1b8, m1r10,m2b 8, m2r 11 (the titles of this set of images are incorrect)
UT 14:46. LUCI1 has a mask in the FPU and the ARGOS cal unit lockout. LUCI2 is blind/blind, mirror and calib out. MODS are asleep and the LBCs are off.

-- BarryRothberg - 29 Aug 2020
Topic revision: r7 - 30 Aug 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback